OpenStudy (anonymous):

solve. 3 1/3 - 2 2/5

OpenStudy (anonymous):

Firstly we look at the denominators: 3 and 5. When faced with these problems, we make the denominators the same. In this case, we have 3 and 5. So logically we will make the denominator 15, since 15 is the lowest common multiple of 3 and 5. Next, we change the numbers. 3 1/3 = 10/3 = 50/15 2 2/5 = 12/5 = 36/150 so 50/15 - 36/15 is 14/15, which I trust that you, too, can calculate. :) Hope this helped. Have a great day! :)

OpenStudy (whpalmer4):

Do you know algebra? If so, there's a nifty trick which makes this much less error-prone than the usual way of doing it. \[3 \frac{1}{3}-2\frac{2}{5} = n\]\(n\) will be our answer Multiply both sides by 3*5=15 to get rid of fractions \[15*(3+\frac{1}{3}) -15*(2+\frac{2}{5}) = 15n\]\[15*3 + 15*\frac{1}{3} - 15*2-15*\frac{2}{5} = 15n\]\[45+5-30-6=15n\]\[14=15n\]\[n=\frac{14}{15}\]